OpenStudy (anonymous):

x^2+6x+8 polynomial expression.

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

example (x+2)(x+5)

OpenStudy (anonymous):

oh, so you need to factor the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay well here's how: x^2+6x+8=0 first you need to find the two factors of eight that will add up to 6 (in this case 4 and 2) (x+2)(x+4) = x^2+6x+8

OpenStudy (anonymous):

y=x^(2)+6x+8 To find the roots)/(zeros of the function, set the function equal to 0 and solve. 0=x^(2)+6x+8 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. x^(2)+6x+8=0 In this problem 4*2=8 and 4+2=6, so insert 4 as the right hand term of one factor and 2 as the right-hand term of the other factor. (x+4)(x+2)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+4=0_x+2=0 Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. x=-4_x+2=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-4_x+2=0 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. x=-4_x=-2 The complete solution is the set of the individual solutions. The multiplicity of a root is the number of times the root appears. For example, a factor of (x+5)^(3) would have a root at x=-5 with multiplicity of 3. x=-4,-2